How many Kilograms of water must be added to 20 kilograms of a 10% salt solution to produce a 5% solution ?
Please help me with not only showing how this is done, but explaining the steps. I am terrible in math!

Answer by guapa(62)

Let x represent the amount of water to be added.
You can use the following guideline:
Amount of water in 10% + Amount of water added = Amount of water in 5% solution
10%(20) + 0%(x) = 5%(20+x)
0.1(20)+ 0 = 0.05(20+x) Now simplify the terms
2+0=1+0.05x
2 = 1 + 0.05x Substract 1 from both sides
2-1 = 1 +0.05x - 1
1 = 0.05x Divide both sides by 0.05
1/0.05 = 0.05x/0.05
x = 20
So, 20kg water must be added.